### v2

# MetaMetrics: Calibrating Metrics For Generation Tasks Using Human Preferences

Abstract
Understanding the quality of a performance evaluation metric is crucial for ensuring that model outputs align with human preferences. However, it remains unclear how well each metric captures the diverse aspects of these preferences, as metrics often excel in one particular area but not across all dimensions. To address this, it is essential to systematically calibrate metrics to specific aspects of human preference, catering to the unique characteristics of each aspect. We introduce MetaMetrics, a calibrated meta-metric designed to evaluate generation tasks across different modalities in a supervised manner. MetaMetrics optimizes the combination of existing metrics to enhance their alignment with human preferences. Our metric demonstrates flexibility and effectiveness in both language and vision downstream tasks, showing significant benefits across various multilingual and multi-domain scenarios. MetaMetrics aligns closely with human preferences and is highly extendable and easily integrable into any application. This makes MetaMetrics a powerful tool for improving the evaluation of generation tasks, ensuring that metrics are more representative of human judgment across diverse contexts.
1 Introduction
Evaluating machine-generated sentences has long been one of the main challenges in natural language processing (NLP). Callison-Burch et al. (2006) provide multiple examples where a high BLEU score does not necessarily indicate true sentence similarity, and conversely, highly similar sentence pairs can receive low BLEU scores. BERTScore (Zhang et al., 2019) is designed to capture semantic similarities, also falls short in capturing the all the nuances to have a comprehensive evaluation. Figure 1 illustrates two such cases where traditional metrics either overestimate or underestimate the alignment of generated sentences. Such cases highlight a critical limitation for robust evaluation, as it may fail to account for the multifaceted aspects of language quality.
In light of recent advancements such as Reinforcement Learning with Human Feedback (RLHF) (Ouyang et al., 2022), ensuring that generated outputs align with human preferences has become increasingly critical. Models that optimize for human preference, rather than solely relying on traditional metrics, have demonstrated superior performance in producing content that aligns with human preference (Rafailov et al., 2024; Winata et al., 2024). This shift highlights the need for evaluation metrics that accurately reflect human subjective judgments across multiple dimensions. Such metrics are essential for guiding models to generate more human-aligned outputs by comparing their results against human judgments. Metrics that exhibit a high correlation with human ratings are considered more reliable and effective for evaluating model performance.
This is particularly important for NLP tasks, where the subtleties of human language and context significantly influence quality assessments. For example, in machine translation (Freitag et al., 2023; Juraska et al., 2023), the accuracy and fluency of the translated text are critical factors considered by humans. Similarly, in text summarization (Fabbri et al., 2021), the coherence, relevance, and conciseness of the summary are key aspects that need evaluation. Beyond NLP tasks, vision-language (VL) tasks like image captioning also benefit from reliable evaluation metrics (Hessel et al., 2021). In these tasks, the generated captions must accurately describe the image content while maintaining grammatical correctness and contextual relevance. The complexity of these tasks highlights the need for metrics that can effectively capture the quality of the generated content in a manner consistent with human judgment. Each metric may be good for a specific human preference aspect, and it is necessary to identify which metric is suitable for each aspect to ensure comprehensive evaluation.
Many Metrics Are Not Tuned on Human Preferences.
Standard metrics that are commonly used to evaluate downstream language tasks, such as BLEU (Papineni et al., 2002) and BERTScore (Zhang et al., 2019), often fail to align with human preferences. In many generation tasks, it is crucial to assess the quality of outputs based on human judgment, as these evaluations comprise multiple dimensions of quality. Therefore, we need a comprehensive framework that allows us to identify, refine, and utilize effective metrics for evaluating generated content in a way that aligns with human judges.
Motivated by this insight, we introduce MetaMetrics, a meta-metric designed to better align with human preferences by calibrating multiple metrics using human assessment scores. Our method is both efficient and fast, eliminating the need for extensive training. Our contributions are three-fold:
  • We introduce MetaMetrics, a supervisedly calibrated, flexible, and modality- and language-agnostic metric that aligns closely with human preferences on five different tasks. Our metric is designed to operate in two distinct settings: reference-based and reference-free, providing versatility across a wide range of tasks and modalities.
  • We demonstrate that MetaMetrics is adaptable to a wide range of requirements, including performance optimization and metric efficiency through a comprehensive benchmarking of diverse metrics across a range of language and vision tasks Section 4. We also show that MetaMetrics can be used effectively as a reward model. [^1] MetaMetrics ranks 4th out of 144 on the overall leaderboard, utilizing the fewest parameters, and secures 2nd place on the detailed leaderboard at RewardBench. The leaderboard reflects data as of October 6th, 2024. This adaptability ensures that MetaMetrics can be calibrated to suit various human preference aspects.
  • We demonstrate that MetaMetrics outperforms existing metrics by leveraging complementary metrics to enhance overall performance. We conduct a detailed analysis to understand how MetaMetrics attributes weight or attention to specific metrics, thereby measuring the contribution of each component. We release the code and models to facilitate reproducibility and empower researchers and practitioners to utilize and extend our metric for various applications. [^2] We release MetaMetrics as an open-source library at https://github.com/meta-metrics/metametrics.
Figure 1: Examples of image captioning on THumB 1.0 dataset comparing MetaMetrics against BLEU and CLIP-S scores. MetaMetrics scores of predicted captions are closer to human ratings compared to BLEU and CLIP-S scores in the left and right images, respectively.
2 Aren’t Existing Metrics Good Enough?
The challenge of developing metrics tailored to specific tasks is not a new issue. For a long time, researchers have struggled to identify appropriate metrics that align with human preferences. In this section, we outline the rationale for the need for a new evaluation metric that is both suitable for our tasks and aligned with human preferences.
Commonly Used Metrics are Not Robust and Unrepresentative.
Metrics such as Perplexity (PPL) (Jelinek et al., 1977; Bengio et al., 2000) and BLEU (Papineni et al., 2002) are widely used to measure the quality of generated text in various generation tasks. However, they are not always the most suitable metrics, particularly when accounting for variations in writing styles and minor character differences due to diacritics (Freitag et al., 2022). Critics have noted that PPL is an English-centric metric that performs well for English but is less effective for other languages, such as Japanese (Kuribayashi et al., 2021). THumB (Kasai et al., 2022) further revealed that widely used metrics fail to adequately capture the semantic quality and diversity of high-quality human annotations, leading to misleading evaluations of model performance, particularly for captions.
Limited Capability of an Individual Metric.
Single metrics like BERTScore (Zhang et al., 2019), ROUGE (Lin, 2004), and METEOR (Banerjee & Lavie, 2005) are beneficial and capable of measuring the quality of generated content. However, they have significant limitations. For instance, in summarization tasks, BERTScore (Recall) excels in assessing consistency but falls short in evaluating coherence. Conversely, BERTScore (F1) performs well in measuring relevance but not consistency (Fabbri et al., 2021). This variability means that a single metric may perform well in one aspect but poorly in another, making it challenging to select the appropriate metric without extensive benchmarking. This ambiguity complicates the process of choosing and utilizing the most suitable metric for a specific use case.
One Metric with Too Many Variants and Implementations.
Metrics like BERTScore allow the use of various BERT models, but the sheer number of options can make it difficult to identify the best one without a systematic approach. This can lead to an exhaustive search process. BERT models trained on English often underperform on non-English languages, and vice versa. Non-English languages in Latin script may also yield poorer results compared to those in their native scripts. Compounding these issues, many metrics are heavily parameterized, yet their settings are frequently undocumented, resulting in score variations across implementations (Post, 2018; Grusky, 2023).
To tackle these challenges, we propose a new paradigm for creating customizable metrics that are robust across various tasks and closely aligned with human preferences. Our approach will enable systematic evaluation and automatic selection of the most suitable combination of metrics that is aligned with human judgments on target tasks, ensuring optimal performance and relevance.
3 MetaMetrics
In this section, we outline the notations and definitions used in our proposed method, and detail the conditions required for optimizing the method based on human preferences. Additionally, we discuss the key factors influencing the optimization process.
Figure 2: MetaMetrics Framework for reference-free ($\theta_{\textsc{MM}}^{\text{noref}}$, left) and reference-based setting ($\theta_{\textsc{MM}}^{\text{ref}}$, right). MetaMetrics integrates multiple metrics $\theta_{i}$ and their scores $\hat{y}_{i}$, learns a function $\Phi$ to combine them into a score $\hat{y}_{\textsc{MM}}$ that aligns best with the human judgment score.
3.1 Preliminaries
Definition.
We define $\theta_{i}$ as a metric function that maps a sample input $x$ to a score $\hat{y}_{i}$, where $i\in\{1,\ldots,N\}$ denotes different metrics. Note that each $\theta_{i}$ depends on the type of task it is applied to. For reference-based metric, the data is evaluated in the context of $x=(x_{\text{hyp}},x_{\text{ref}})$, where $x_{\text{hyp}}$ and $x_{\text{ref}}$ correspond to the hypothesis text and the reference text, respectively. For reference-free metric, only $x_{\text{hyp}}$ will be used. In VL tasks, the input is extended to include an image, which would correspond to $x=(x_{\text{text}},x_{\text{image}})$ where $x_{\text{text}}$ and $x_{\text{image}}$ correspond to the caption and image, respectively.
Given a set of $N$ evaluation metrics $\{\hat{y}_{1},\ldots,\hat{y}_{N}\}$, we define $\Phi$ to compute a scalar meta-metric score of $\hat{y}_{\textsc{MM}}$. The idea of utilizing multiple metrics is to combine scores from multiple metrics regardless of the metric types. Overall, we define $\theta_{\textsc{MM}}$ as a meta-metric function where $\hat{y}_{\textsc{MM}}$ is computed as follows:
$\hat{y}_{i}=\theta_{i}(x),\quad\hat{y}_{\textsc{MM}}=\theta_{\textsc{MM}}(x)=\Phi(\hat{y}_{1},\cdots,\hat{y}_{N}).$ (1)
The MetaMetrics $\theta_{\textsc{MM}}$ is used to calculate the objective value $\rho(\hat{y}_{\textsc{MM}},z)$, where $\rho$ is measures alignment with $z\in\mathbb{R}$, the human preference score. In practice, $z$ can represent various scores annotated by human judges, such as coherence and fluency.
3.2 Human Preference Optimization
Notations and Objective.
Recall that we aim to calibrate $\theta_{\textsc{MM}}$, that maximizes an objective calibration function $\rho(\hat{y}_{\text{MM}},z)$, where $z$ denotes human assessment scores–encompassing any score annotated by human evaluators. $\rho$ is a function that measures the alignment between $z$ and MetaMetrics scores, $\hat{y}_{\text{MM}}$. MetaMetrics is designed to combine scores of multiple metrics $\theta_{1}(x),\theta_{2}(x),\ldots,\theta_{N}(x)$, learning the weights $w_{i}$ to assign to each $\hat{y}_{i}=\theta_{i}(x)$ to maximize $\rho(\hat{y}_{\text{MM}},z)$. Each metric has its score $\hat{y}_{i}$ ranges within a specific minimum and maximum value. Some metrics, particularly neural-based ones, can fall outside this defined range. To standardize these metrics, we need to normalize them to a common scale that ranges from 0 to 1. In this scale, 0 represents poor translation performance, while 1 represents perfect translation performance. We apply pre-processing to $\hat{y}_{i}$ before we learn to combine these scores in MetaMetrics training. The pre-processing is defined in Section B.2.
The advantage of MetaMetrics lies in its flexibility and adaptability to different tasks or domains. Certain metrics on certain tasks may exhibit strong correlations with human judgments, and by constructing a composite metric that learns to integrate these reliable metrics with human judgments, we can enhance the overall correlation with human evaluations.
3.3 Optimization Methods
In this work, we focus on two optimization methodologies to train MetaMetrics: Bayesian Optimization (BO) and Boosting. BO offers the advantage of interpretability, allowing us to clearly identify which metrics contribute most significantly to the final outcome. Conversely, the Boosting method excels in enhancing alignment and accounting for the compositionality of different metrics, even when dealing with more complex functions. Although we can measure the contribution of each metric, the clarity and distinctness of these contributions are more pronounced with BO compared to Boosting.
3.3.1 Bayesian Optimization (BO)
BO constructs a posterior distribution of functions, typically modeled as a Gaussian Process (GP), to represent the function being optimized. As observations accumulate, the posterior distribution becomes more precise, allowing the algorithm to identify which regions of the parameter space to explore and which to ignore. BO is particularly effective for predicting unknown functions based on observed data, making it a robust method for optimizing complex and costly-to-evaluate functions such as the correlation between evaluation metric scores and human assessment scores.
Formally, we want to solve for the optimum weights $\mathbf{w}^{*}=\arg\max_{\mathbf{w}\in\mathbf{W}}\rho(\hat{y}_{\textsc{MM}}(\mathbf{w}),z)$, where $\rho(\hat{y}_{\textsc{MM}}(\mathbf{w}),z)$ is the alignment between MetaMetrics score $\hat{y}_{\textsc{MM}}(\mathbf{w})$ with the human preference score $z$, $\mathbf{w}=[w_{1},w_{2},\ldots,w_{N}]$, $\hat{y}_{\textsc{MM}}(\mathbf{w})=w_{1}\hat{y}_{1}+w_{2}\hat{y}_{2}+\ldots+w_{N}\hat{y}_{N}$, and $\mathbf{W}=[\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{K}]$ is the set of weight vectors.
We assume a GP prior: that the alignment between the weighted combination of metric scores and the human preference score follows a GP $\rho(\hat{y}_{\textsc{MM}}(\mathbf{w}),z)\sim\mathcal{GP}(\mu(\mathbf{w}),k(\mathbf{w},\mathbf{w^{\prime}}))$, where $\mu(\mathbf{w})$ is is the prior mean of the alignment and $k(\mathbf{w},\mathbf{w^{\prime}})$ is the kernel function that measures the similarity between two sets of weights $\mathbf{w}$ and $\mathbf{w^{\prime}}$. After observing a set of weight vectors $\mathbf{W}=[\mathbf{w}_{1},\mathbf{w}_{2},\dots,\mathbf{w}_{k}]$ and their corresponding alignments $\mathbf{\rho}=[\rho_{1},\rho_{2},\dots,\rho_{k}]$, the GP posterior can be updated to predict the alignment for a new weight vector $\mathbf{w}$. To guide the search for the optimal weights $\mathbf{w}$, BO maximizes a function and selects the next weight vector $\mathbf{w}_{k+1}$ to evaluate using the function. BO’s function balances exploration and exploitation using the posterior mean and variance, iterating until convergence or budget exhaustion. While challenging, maximizing the function is computationally efficient with standard tools, making BO suitable for resource-intensive sampling scenarios. GP employs a multivariate Gaussian distribution, offering a sparse representation of metrics that enhances interpretability by identifying key contributors to the objective.
3.3.2 Boosting Method
The boosting method we investigate is Extreme Gradient Boosting (XGBoost) (Chen & Guestrin, 2016). Gradient-boosted trees have long been recognized as a robust technique, supported by extensive literature demonstrating their effectiveness (Friedman, 2001). To enhance the efficiency of our metric, we implement iterative pruning to eliminate less important metrics from the input, resulting in a more compact and faster metric.
Metric SummEval Benchmark LLM Combined
Coh. Cons. Fluency Rel. Faith. Coh. Cons. Coh. Cons. Avg.
BLEU 0.110 0.126 0.113 0.170 0.162 0.246 0.252 0.029 0.082 0.143
chrF 0.143 0.094 0.071 0.198 0.272 0.163 0.342 0.068 0.218 0.174
METEOR 0.077 0.102 0.072 0.162 0.251 0.182 0.331 0.033 0.164 0.153
ROUGE-WE1 0.115 0.088 0.081 0.169 0.223 0.218 0.292 0.198 0.272 0.184
BARTScore (Mean) 0.086 0.074 0.040 0.143 0.080 0.216 0.166 0.210 0.173 0.132
BLEURT (Max) 0.185 0.070 0.114 0.189 0.046 0.264 0.143 0.229 0.143 0.154
BERTScore (f1) 0.105 0.100 0.120 0.181 0.171 0.280 0.283 0.306 0.304 0.206
MetaMetrics-Sum
XGBoost 0.190 0.151 0.131 0.271 0.406 0.411 0.397 0.403 0.367 0.303
XGBoost (Iterative Top 5) 0.141 0.115 0.160 0.208 0.377 0.378 0.383 0.369 0.351 0.276
GP 0.170 0.142 0.115 0.233 0.356 0.308 0.383 0.315 0.333 0.262
Table 1: Kendall correlation results with human ratings for summarization task. Coh., Cons., Rel., and Faith. corresponds to coherence, consistency, relevance, and faithfulness respectively. We only present the best-performing metric variant from each metric category, more detailed results for all metrics and their variants can be found in Table 9 in the Appendix. Bold and underlined values indicate the best and second best performance, respectively. Our MetaMetrics-SUM correlates the best with human ratings across different datasets and aspects of evaluation.
Iterative-based Pruning.
When training an XGBoost regressor model, we can initially use a large set of metrics. However, this approach is less efficient than the BO model at inference time. To address the scalability and efficiency issues, we propose an iterative pruning method (Algorithm 1). We conduct $k$ iterations of XGBoost training, starting from the full set of metrics and removing the metric with the least feature importance at each iteration. We return the best XGBoost model among $k$ iterations based on the cross-validation value of $\rho(\hat{y}_{\text{MM}},z)$ measured during training.
4 Experiments Setup
In this work, we explore two optimization methodologies: BO and Boosting. BO provides interpretability, clearly highlighting the metrics that significantly impact the final outcome. In contrast, Boosting enhances alignment and addresses the compositionality of metrics, even for complex functions. While we can measure each metric’s contribution through feature importance in Boosting, BO is more interpretable than Boosting. We also compare calibrating with all the metrics and calibrating only using top-5 correlated metrics from the tuning set.
Abstractive Text Summarization.
We use the SummEval (Fabbri et al., 2021) and BenchmarkLLM (Zhang et al., 2024) datasets for text summarization evaluation. SummEval, based on CNN/DailyMail (Hermann et al., 2015), is rated by human annotators on coherence, consistency, fluency, and relevance using a 1 to 5 Likert scale. BenchmarkLLM includes both CNN/DailyMail and XSUM (Narayan et al., 2018) articles, with summaries assessed for faithfulness (binary scale), coherence, and consistency (1 to 5 Likert scale). To evaluate the generalization capabilities of MetaMetrics, we combine coherence and consistency scores from both datasets, ensuring no overlap in articles, allowing us to assess MetaMetrics’s performance across different human annotators. We use the Kendall $\tau$ correlation function as our objective calibration function $\rho$. We refer to our metric as MetaMetrics-SUM (Table 1).
Machine Translation.
For this task, we train both reference-free and reference-based versions of the metric. We train our metric using the 3-year training data from WMT shared tasks datasets from 2020 to 2022 that are annotated using MQM annotation scores. We evaluate on MQM dataset from WMT23 shared task. We use kendall $\tau$ correlation function as our objective calibration function $\rho$. We refer to our metric as MetaMetrics-MT (Table 2).
Metric overall en-de he-en zh-en
sys/seg sys seg seg sys seg seg sys seg seg
avg-corr pearson pearson acc-t pearson pearson acc-t pearson pearson acc-t
Reference-based Metric
    chrF 0.694 0.866 0.232 0.519 0.776 0.221 0.460 0.809 0.063 0.485
    BLEU 0.696 0.917 0.192 0.520 0.769 0.220 0.442 0.734 0.119 0.472
    BERTScore 0.742 0.891 0.325 0.528 0.895 0.335 0.515 0.810 0.236 0.499
    Yisi-1 0.754 0.925 0.366 0.542 0.917 0.395 0.529 0.823 0.290 0.504
    MetricX-23-XXL 0.808 0.977 0.585 0.603 0.910 0.548 0.577 0.873 0.625 0.531
    XCOMET-Ensemble 0.825 0.980 0.695 0.604 0.950 0.556 0.586 0.927 0.650 0.543
    COMET 0.779 0.990 0.432 0.575 0.940 0.401 0.531 0.898 0.396 0.514
MetaMetrics-MT
    GP 0.819 0.970 0.638 0.610 0.947 0.546 0.590 0.900 0.646 0.539
    XGBoost 0.825 0.992 0.680 0.616 0.957 0.557 0.574 0.929 0.637 0.546
Reference-free Metric
   GEMBA-MQM 0.802 0.993 0.502 0.572 0.939 0.401 0.564 0.991 0.449 0.522
   XCOMET-QE-Ensemble 0.808 0.974 0.679 0.588 0.909 0.498 0.554 0.892 0.647 0.533
   MetricX-23-QE-XXL 0.797 0.934 0.547 0.607 0.813 0.459 0.575 0.877 0.652 0.531
   CometKiwi-QE-XL 0.786 0.976 0.447 0.571 0.900 0.384 0.533 0.974 0.430 0.522
MetaMetrics-MT-QE
   GP 0.801 0.934 0.556 0.609 0.815 0.474 0.578 0.900 0.660 0.537
   XGBoost 0.805 0.967 0.583 0.604 0.881 0.509 0.568 0.869 0.642 0.526
Table 2: Kendall correlation results with human ratings on WMT23 (MQM). The results are collected from Freitag et al. (2023). More detailed results for all metrics and their variants can be found in Table 11 in the Appendix. Bold and underlined values indicate the best and second best performance, respectively. Our MetaMetrics-MT correlates with human judgments better than or comparably to the best WMT evaluation metrics across different languages and evaluation settings.
Question Answering.
For this task, we evaluate MetaMetrics performance across multiple QA subtasks, including Open Domain QA (Open-QA) (Rajpurkar, 2016; Berant et al., 2013; Petroni et al., 2019), Reading Comprehension QA (RCQA) (Bajaj et al., 2016; Fisch et al., 2019; Yang et al., 2018), and Reasoning QA. Open-QA retrieves answers from large, unstructured datasets like Wikipedia, where answers are objective. RCQA requires the model to read a passage to derive answers directly from it, while Reasoning QA emphasizes logical inference, where answers cannot be directly extracted from the passage.
Due to the nature of our paper, we require QA datasets with human evaluations. The data we use are collected from two sources: Evaluating Question Answering Evaluation (EQAE) (Chen et al., 2019) and EVOUNA (Wang et al., 2024a) and include Open-QA datasets: Natural Questions (NQ) (Kwiatkowski et al., 2019) and TriviaQA (TQ) (Joshi et al., 2017), RCQA dataset: NarrativeQA (Kočiskỳ et al., 2018), and Reasoning QA dataset: SemEval 2018 Task 11 (SemEval) (Ostermann et al., 2018). We refer to our metric as MetaMetrics-QA (Table 3).
Setup EVOUNA EQAE Combined
NQ TQ NarrativeQA SemEval
ArmoRM Metrics
ultrafeedback-honesty 0.202 0.281 0.297 0.335 0.296
helpsteer-helpfulness 0.204 0.293 0.332 0.427 0.302
helpsteer-correctness 0.206 0.300 0.336 0.403 0.305
argilla-judge-lm 0.243 0.301 0.253 0.337 0.332
n-gram-based Metrics
BLEU1 0.461 0.353 0.362 0.260 0.424
ROUGE-L 0.495 0.399 0.584 0.518 0.454
METEOR 0.517 0.435 0.570 0.461 0.494
MetaMetrics-QA
GP 0.518 0.448 0.586 0.464 0.512
XGBoost 0.543 0.471 0.601 0.486 0.536
XGBoost (Iterative Top 5) 0.540 0.488 0.626 0.503 0.528
Table 3: Kendall correlation results of MetaMetrics-QA. More detailed results for all metrics and their variants can be found in Table 10 in the Appendix. Bold and underlined values indicate the best and second best performance, respectively. In most cases, our MetaMetrics-QA outperforms the other metrics in correlation with human judgments.
Image Captioning.
We use Flickr8k-Expert (Hodosh et al., 2013) and THumB 1.0 (Kasai et al., 2022) human evaluation datasets. Flickr8k-Expert consists 5,882 image-caption pairs, each having 3 human annotators with a scale of 1–4 where 1 is worst and 4 is best. The THumB 1.0 dataset is a rubric-based human evaluation framework applied to a subset of the MSCOCO dataset (Lin et al., 2014). The final score is of a scale of 1–5 where 1 is worst and 5 is best. We refer to our metric as MetaMetrics-Cap (Figure 3).
Reward Model Scoring.
We use MetaMetrics as reward model to score text generation from LLM. We use RewardBench (Lambert et al., 2024) as the test benchmark. We use Skywork Reward Data Collection [^3] Skyword Reward Data Collection can be accessed at https://huggingface.co/datasets/Skywork/Skywork-Reward-Preference-80K-v0.1 for training our MetaMetrics, including HelpSteer2 (Wang et al., 2024c), OffsetBias (Park et al., 2024), WildGuard (Han et al., 2024), and Magpie (Xu et al., 2024). In addition, we add Preference Test Data [^4] Preference Test Data from AI2 can be accessed at https://huggingface.co/datasets/allenai/preference-test-sets to increase the size of the human preference dataset. We refer to our metric as MetaMetrics-RM and report the accuracy of predicting that the chosen sample is ranked higher than the rejected sample (Table 4).
5 Results
In this section, we present the results of our MetaMetrics across five diverse tasks, covering both NLP and VL downstream applications.
Abstractive Text Summarization.
For this task, we refer to our metric as MetaMetrics-SUM. Table 9 shows the results of abstractive text summarization task. Overall, our proposed model with XGBoost outperforms all other baselines, including all ensemble models and overall best automatic metric (BERTScore (F1)). An interesting observation we find is our iterative model with fewer features achieves slightly better average correlation score ($\Delta$ 0.004) compared to the model with full metrics. The MetaMetrics-SUM with GP also performs considerably well, also outperforms all other baselines.
Figure 3: Kendall correlation results of MetaMetrics-Cap. Cross-Dataset means that MetaMetrics is tuned on THumB 1.0 and tested on Flickr8k, and vice versa. Our MetaMetrics-Cap outperforms other metrics across different datasets and shows robustness even when trained and tested on different datasets (Cross-Dataset setting). More detailed results for all metrics and their variants can be found in Table 12 in the Appendix.
Metric Score Chat Chat Hard Safety Reasoning
Llama-3.1-Nemotron-70B-Reward 94.1 97.5 85.7 95.1 98.1
Skywork-Reward-Gemma-2-27B 93.8 95.8 91.4 91.9 96.1
TextEval-Llama3.1-70B 93.5 94.1 90.1 93.2 96.4
Skywork-Critic-Llama-3.1-70B 93.3 96.6 87.9 93.1 95.5
SFR-LLaMa-3.1-70B-Judge-r 92.7 96.9 84.8 91.6 97.6
URM-LLaMa-3.1-8B 92.9 95.5 88.2 91.1 97.0
Skywork-Reward-Llama-3.1-8B 92.5 95.8 87.3 90.8 96.2
GRM-Llama3-8B 91.5 95.5 86.2 90.8 93.6
ArmoRM-Llama3-8B-v0.1 90.4 96.9 76.8 90.5 97.3
URM-LLaMa-3-8B 89.9 96.9 78.7 88.2 95.7
MetaMetrics-RM (GP) 93.4 98.3 86.4 90.8 98.2
MetaMetrics-RM-XGBoost
   Iterative Best 92.9 95.8 88.4 91.8 95.8
Table 4: Accuracy results in percentages for Reward-Model-As-A-Metric. More detailed results for all metrics and their variants can be found in Table 13 in the Appendix. Values that are bolded indicate the best performance, while those that are underlined represent the second-best performance. Our MetaMetrics-RM achieves the best or comparable results in 2 out of the 4 categories in RewardBench.
Machine Translation.
In this task, we demonstrate that MetaMetrics-MT in the reference-based setting outperforms all baseline models, including ensemble methods (Table 2). Notably, the top competitor, XCOMET-Ensemble, achieves impressive results on the WMT23 evaluation leaderboard, particularly excelling in the two-segment Pearson correlation metric for en-de and zh-en. However, MetaMetrics consistently surpasses XCOMET-Ensemble across all other tasks, leading to a higher average correlation overall, achieving the new state-of-the-art performance on WMT23 evaluation.
In the reference-free setting, MetaMetrics-MT-QE performs comparably to XCOMET-QE-Ensemble. While XCOMET-Ensemble shows slightly better average correlation, our proposed model demonstrates superior performance on the acc-t metric. This is particularly significant, as the acc-t metric is finely tuned to align with the Kendall correlation $\tau$ objective function, indicating a stronger correlation with human evaluations.
Question Answering.
Table 3 presents the results of MetaMetrics-QA, which consistently outperforms both standard and existing metrics across all evaluation levels. While ROUGE performs slightly better than our method in the SemEval dataset, our proposed model demonstrates a significant advantage over others in different datasets, achieving improvements of up to 0.044 points. Notably, our iterative-based approach enhances the performance of the vanilla MetaMetrics-QA when using XGB, all while utilizing fewer metrics. Additionally, for the GP evaluation, selecting all metrics or just the top five yields comparable combined scores, indicating the robustness of our metric selection strategy.
Image Captioning.
Figure 3 presents the results of MetaMetrics-Cap and cross-dataset tuning. MetaMetrics-Cap consistently performs better compared to any individual metric, even when the tuning is performed on a dataset from a different distribution. Although tuning on the same distribution yields the best results, we observe that MetaMetrics-Cap still outperforms any individual metric, demonstrating its robustness across varying data distributions.
Metric Model Scoring.
Table 4 highlights the performance of MetaMetrics-RM, which achieves state-of-the-art results [^5] We use the RewardBench leaderboard from October 6th, 2024, as the basis for comparison. for GP and competes effectively with the top seven metrics, including the advanced 70B Nemotron Reward model. These results demonstrate that our proposed metric is both efficient and effective, showcasing its robustness across diverse categories of tasks, datasets, and evaluation settings. Overall, MetaMetrics-RM represents a significant advancement in reward metric evaluation, offering high performance and broad applicability for researchers and practitioners alike.
6 Analysis and Discussion
In this section, we analyze the advantages of the methods in terms of interpretability, efficiency, and robustness. Additionally, we discuss how MetaMetrics enhances the capability of model evaluation.
6.1 Interpretability
Figure 4: GP weights (Left) and XGBoost Features Importance (Center) with Intra-Metric Correlation (Right) for MetaMetrics-QA.
Interpretability is a key aspect in MetaMetrics, as the optimization process inherently reveals the impact of each metric on the final score. The chosen methods, such as feature importance in Boosting and weights analysis in Bayesian Optimization, provide a straightforward way to quantify each metric’s contribution to human preference alignment. As shown in Figure 4, this interpretability is highlighted in the QA task, showcasing the importance of each metric in human preference alignment. For example, since ROUGE and ROUGE-L are two highly correlated metrics, Bayesian Optimization often drops one of them during the tuning process.
6.2 Efficiency
Improved Performance with Fewer Parameters.
Table 4 highlights the compute efficiency and superior performance of MetaMetrics-RM. MetaMetrics-RM, utilizing GP, combines GRM-Gemma-2B, Internlm2-7B-Reward, Skywork-Reward-Llama-3.1-8B, Internlm2-1.8B-Reward, and GRM-Llama3-8B totaling approximately 27B parameters. Despite this relatively small compute memory footprint, it outperforms models more than 2$\times$ its size. This demonstrates that MetaMetrics-RM offers state-of-the-art accuracy using significantly more cost-effective models, making it an ideal choice for resource-constrained applications while delivering even better performance.
High Parallelization.
MetaMetrics is designed to be embarrassingly parallel, enabling each metric to be executed independently without requiring inter-model communication, which results in exceptional efficiency. Unlike large models with 27 billion parameters that often face bottlenecks, our approach supports the development of highly effective models without such time constraints. Furthermore, MetaMetrics excels at selecting a sparse yet impactful set of metrics from a vast pool of potential evaluation metrics. As shown in Figure 4, both GP and XGBoost can identify a small subset of key metrics that strongly correlate with human preferences, thereby reducing the need to evaluate a wide range of metrics. This selective approach identifies essential metrics for specific tasks while minimizing computational overhead without compromising performance. In essence, MetaMetrics is ideally suited for low-resource settings, providing superior alignment with human preferences across multiple tasks.
6.3 Robustness
We evaluate the robustness of MetaMetrics through cross-dataset experiments in image captioning, tuning on one dataset and testing on another. Despite variations in domain and content, our results consistently demonstrate the effectiveness of MetaMetrics. Figure 3 shows that MetaMetrics-Cap tuned on THumB 1.0 and tested on Flickr8k and vice-versa out performs all individual metrics. This robustness indicates that MetaMetrics is not overfitting to specific dataset characteristics, but rather learning to combine metrics in a way that aligns with general human judgments of caption quality. This generalization capability is crucial for real-world applications, where the evaluation metric may need to perform well on diverse and potentially out-of-domain data.
7 Related Work
Performance evaluation metrics for natural language generation tasks can be categorized into several categories based on how the metric compares generated texts with reference texts.
Surface-level Metrics.
These metrics compare the generated text to reference text at the word level, focusing on n-gram overlap or direct lexical matching. Common metrics include BLEU that measures precision of n-grams between the generated- and reference-texts (Papineni et al., 2002), ROUGE that measures recall (Lin, 2004), METEOR (Banerjee & Lavie, 2005) that goes beyond exact word overlap and considers stemming, synonyms, and paraphrasing, and chrF (Popović, 2015) that calculates similarity based on character n-grams rather than word n-grams using F-score.
Embedding-based Metrics.
These metrics rely on word or sentence embeddings to measure semantic similarity between generated and reference text, capturing deeper meaning rather than surface overlap. Examples include MoverScore (Zhao et al., 2019), which considers semantic similarity at the word level and accounts for word movement (alignment) between texts, and BERTScore (Zhang et al., 2019), a neural-based metric that compares semantic similarity and understands contextual relationships using BERT embeddings. COMET (Rei et al., 2022) and BLEURT (Sellam et al., 2020) are neural-based metrics that use contextualized embeddings from models like BERT as inputs to train models to generate prediction estimates of human judgments. For vision-language tasks, ClipScore (Hessel et al., 2021) uses embeddings to compare vision and language modalities.
Ensemble-based Metrics.
Ensemble-based metrics combine multiple individual metrics to enhance evaluation robustness through averaging (Adiwardana et al., 2020) or heuristics (Phy et al., 2020). The Absolute-Rating Multi-Objective Reward Model (ArmoRM) (Wang et al., 2024b) is a SotA reward model designed for RLHF (Ouyang et al., 2022). ArmoRM uses absolute ratings across 19 objectives trained to predict human preference scores.
8 Conclusion
Understanding the quality of a performance evaluation metric is crucial for ensuring alignment with human preferences. However, it remains unclear how well each metric captures the diverse aspects of these preferences, as metrics often excel in one particular area but not across all dimensions. To address this, it is essential to systematically calibrate metrics to specific aspects of human preference, catering to the unique characteristics of each aspect. We introduce MetaMetrics, a calibrated meta-metric designed to evaluate generation tasks across different modalities in a supervised manner. MetaMetrics optimizes the combination of existing metrics to enhance their alignment with human preferences. Our method demonstrates flexibility and effectiveness in both language and vision downstream tasks, showing significant benefits across various multilingual and multi-domain scenarios. Our proposed metric aligns closely with human preferences and it is also highly extendable and easily integrable into any application. This makes it a powerful tool for improving the evaluation of generation tasks, ensuring that metrics are more representative of human judgment across diverse contexts.
Limitations and Future Works
In our experiment, we utilize metrics that are commonly employed to evaluate each downstream task. However, due to limitations in computational resources, we exclude LLMs exceeding 10.7B parameters to ensure that our metrics can be executed on commercial GPU resources with a maximum memory capacity of 40GB. Additionally, we restrict our exploration of metrics to avoid exhaustively incorporating LLM-based metrics, given our capacity and resource constraints.
Looking ahead, there are significant opportunities for further exploration. We can investigate metrics across additional languages and in larger multilingual generation tasks, where we identify MetaMetrics for multilingual applications (Winata et al., 2019). Furthermore, exploring metrics in the context of speech modalities presents another promising avenue for research.
Appendix A Metrics
In this section, we list all metrics we use in our experiments.
A.1 Abstractive Text Summarization
We use the following metrics:
n-Gram-based Metrics: BLEU-4 (Papineni et al., 2002), chrF (Popović, 2015), METEOR (Banerjee & Lavie, 2005), ROUGE-1, ROUGE-3 , ROUGE-4, ROUGE-L (Lin, 2004), ROUGE-WE1 (Ng & Abrecht, 2015).
Embedding-based Metrics: BARTScore (Yuan et al., 2021), BLEURT (Sellam et al., 2020), BERTScore (Zhang et al., 2019).
LLM-based Metrics: SummaQA (Scialom et al., 2019).
A.2 Machine Translation
We use the following metrics:
n-Gram-based Metrics: chrF (Popović, 2015), BLEU (Papineni et al., 2002);
Embedding-based Metrics: BERTScore (Zhang et al., 2019), Yisi-1 (Lo, 2020);
LLM-based Metrics: MetricX-23 (L, XL, XXL) (Juraska et al., 2023), COMET, CometKiwi (Rei et al., 2022), XCOMET (Guerreiro et al., 2023), GEMBA-MQM (Kocmi & Federmann, 2023).
For machine translation, we optimize the Kendall correlation during our tuning.
A.3 Question Answering
We use the following metrics:
n-Gram-based Metrics: BLEU-1, BLEU-4 BLEU (Papineni et al., 2002), ROUGE, ROUGE-L (Lin, 2004), METEOR (Banerjee & Lavie, 2005);
Embedding-based Metrics: BERTScore (Zhang et al., 2019);
LLM-based Metrics: ArmoRM-Llama3-8B-v0.1 (ArmoRM) (Wang et al., 2024b).
A.4 Image Captioning
We use the following metrics:
n-Gram based metrics BLEU (Papineni et al., 2002), ROUGE (Lin, 2004), METEOR (Banerjee & Lavie, 2005), SPICE (Anderson et al., 2016), CIDEr (Vedantam et al., 2015);
Embedding-based Metrics TIGEr (Jiang et al., 2019) and CLIPScore (Both reference-based and reference-free) (Hessel et al., 2021).
A.5 Reward Model Scoring
The reward model used in the experiment is listed in Table 5.
Reward Model / Metric Hugging Face Link
nvidia/Llama-3.1-Nemotron-70B-Reward https://huggingface.co/nvidia/Llama-3.1-Nemotron-70B-Reward
SF-Foundation/TextEval-Llama3.1-70B https://huggingface.co/SF-Foundation/TextEval-Llama3.1-70B
Skywork/Skywork-Critic-Llama-3.1-70B https://huggingface.co/Skywork/Skywork-Critic-Llama-3.1-70B
Salesforce/SFR-LLaMa-3.1-70B-Judge-r https://huggingface.co/Salesforce/SFR-LLaMa-3.1-70B-Judge-r
internlm/internlm2-1_8b-reward (Cai et al., 2024) https://huggingface.co/internlm/internlm2-1_8b-reward
internlm/internlm2-7b-reward (Cai et al., 2024) https://huggingface.co/internlm/internlm2-7b-reward
LxzGordon/URM-LLaMa-3-8B https://huggingface.co/LxzGordon/URM-LLaMa-3-8B
LxzGordon/URM-LLaMa-3.1-8B https://huggingface.co/LxzGordon/URM-LLaMa-3.1-8B
Ray2333/GRM-Gemma-2B-rewardmodel-ft https://huggingface.co/Ray2333/GRM-Gemma-2B-rewardmodel-ft
Ray2333/GRM-Llama3-8B-rewardmodel-ft https://huggingface.co/Ray2333/GRM-Llama3-8B-rewardmodel-ft
sfairXC/FsfairX-LLaMA3-RM-v0.1 https://huggingface.co/sfairXC/FsfairX-LLaMA3-RM-v0.1
Skywork/Skywork-Reward-Llama-3.1-8B https://huggingface.co/Skywork/Skywork-Reward-Llama-3.1-8B
weqweasdas/RM-Mistral-7B https://huggingface.co/weqweasdas/RM-Mistral-7B
Table 5: Reward Models.
Appendix B Datasets
In this section, we present the data statistics and outline the preprocessing procedures used in our work.
B.1 Data Statistics
Table 6 presents the dataset split sizes for tuning and testing.
Dataset Name Type Tuning Size Testing Size
Machine Translation
WMT23 (MQM) (Freitag et al., 2023) Translation - 136,260
WMT20-22 (MQM) (Freitag et al., 2022) Translation 150,343 -
Text Summarization
SummEval (Fabbri et al., 2021) Summarization 510 1,190
BenchmarkLLM (Zhang et al., 2024) Summarization 1,035 2,415
Combined Summarization 1,545 3,605
Question Answering
NQ (Kwiatkowski et al., 2019) Open-QA 4,528 10,567
TQ (Joshi et al., 2017) Open-QA 2,906 6,783
NarrativeQA (Kočiskỳ et al., 2018) RCQA 150 350
SemEval (Ostermann et al., 2018) Reasoning QA 90 210
Combined Multidomain-QA 7,675 17,909
Image Captioning
Flickr8k-Expert (Hodosh et al., 2013) Captioning 1,746 4,076
THumB 1.0 (Kasai et al., 2022) Captioning 600 1,400
Reward Model Scoring
Skywork-Reward-Preference-80K-v0.1 Math, Code, Chat 81,973 -
allenai/preference-test-sets Math, Code, Chat 43,175 -
RewardBench (Lambert et al., 2024) Chat, Safety, Reasoning - 8,108
Table 6: Dataset statistics used in the experiments.
B.2 Data Preprocessing
Pre-processing.
We apply pre-processing before we use the scores in the training. The pre-processing can be defined as follows:
  1. 1.
    Clipping: Let the valid range for $\hat{y}_{i}$ be defined by $[\hat{y}_{i}^{\text{min}},\hat{y}_{i}^{\text{max}}]$. The clipped metric $\hat{y}_{i}^{\prime}$ can be defined as:
    $\hat{y}_{i}^{\prime}=\begin{cases}\hat{y}_{i}^{\text{min}}&\text{if }\hat{y}_{i}<\hat{y}_{i}^{\text{min}},\\ \hat{y}_{i}&\text{if }\hat{y}_{i}^{\text{min}}\leq\hat{y}_{i}\leq\hat{y}_{i}^{\text{max}},\\ \hat{y}_{i}^{\text{max}}&\text{if }\hat{y}_{i}>\hat{y}_{i}^{\text{max}}.\end{cases}$ (2)
  2. 2.
    Normalization: After clipping, the score is normalized to a common scale of $[0,1]$:
    $\hat{y}_{i}=\frac{\hat{y}_{i}^{\prime}-\hat{y}_{i}^{\text{min}}}{\hat{y}_{i}^{\text{max}}-\hat{y}_{i}^{\text{min}}}.$ (3)
  3. 3.
    Inversion (if applicable): If the metric is such that higher scores indicate worse performance, we invert the normalized score:
    $\hat{y}_{i}=1-\hat{y}_{i}.$ (4)
Appendix C Optimization
In this section, we explore the optimization details of MetaMetrics.
C.1 Method Details
C.1.1 Matern Kernel
We train BO using GP with Matern kernel, a generalization of the RBF, distinguished by an additional parameter that controls the smoothness of the resulting function (Williams & Rasmussen, 2006). Conversely, as the parameter approaches infinity, the Matern kernel converges to the RBF kernel. The kernel is described as below:
$k(\mathbf{w},\mathbf{w^{\prime}})=\frac{1}{\Gamma(\nu)2^{\nu-1}}\Bigg{(}\frac{\sqrt{2\nu}}{l}d(\mathbf{w},\mathbf{w^{\prime}})\Bigg{)}^{\nu}K_{\nu}\Bigg{(}\frac{\sqrt{2\nu}}{l}d(\mathbf{w},\mathbf{w^{\prime}})\Bigg{)},$ (5)
where $d(\cdot,\cdot)$ is the Euclidean distance, $K_{\nu}(\cdot)$ is a modified Bessel function and $\Gamma(\cdot)$ is the gamma function.
C.1.2 Iterative-based Pruning
Below is the implementation of Iterative-based Pruning performed during training.
Algorithm 1 Iterative-based Pruning with XGBoost
1:procedure IterativeXGBoost($\mathbf{X},\mathbf{y},k$)
2:     $\mathcal{F}\leftarrow{\{f_{1},f_{2},\dots,f_{p}\}}$triangleright Initial set of features
3:     $\mathcal{P}\leftarrow[]$triangleright Performance history
4:     $\mathcal{F_{\text{least}}}\leftarrow[]$triangleright Feature pruning history
5:     for $i\leftarrow 1$ to $k$ do
6:         Train $\Phi^{(i)}_{\textsc{XGB}}$ on $\mathbf{X}_{\mathcal{F}}$ with CV
7:         $\mathcal{I}_{i}\leftarrow\text{Importance}(\Phi_{\textsc{XGB}})$triangleright Compute feature importance from $\Phi_{\textsc{XGB}}$
8:         $\mathcal{P}[i]\leftarrow\rho(\Phi_{\textsc{XGB}}(\mathbf{X}_{\mathcal{F}}),\mathbf{y})$triangleright Store performance score
9:         $f_{\text{least}}[i]\leftarrow\text{argmin}(\mathcal{I}_{i})$triangleright Identify least important feature
10:         $\mathcal{F_{\text{least}}}\leftarrow f_{\text{least}}$
11:         $\mathcal{F}\leftarrow\mathcal{F}\setminus\{f_{\text{least}}\}$triangleright Remove least important feature
12:     end for
13:     $i^{*}\leftarrow\text{argmax}(\mathcal{P})$triangleright Find the best index iteration
14:     $\mathcal{F}_{\text{best}}\leftarrow\mathcal{F}\cup\mathcal{F}_{\text{least}}[i^{*}:]$triangleright Best features from highest performance
15:     Train final $\hat{f}_{\textsc{XGB}}$ on $\mathbf{X}_{\mathcal{F}_{\text{best}}}$
16:     return $\hat{f}_{\text{xgb}}$
17:end procedure
C.2 Hyper-parameters
C.2.1 Bayesian Optimization
Table 7 describes the hyper-parameter settings that we use for our experiments. For the Bayesian optimization, we run GP with a Matern kernel Williams & Rasmussen (2006), a generalization of the RBF kernel, using $\nu=2.5$.
Metric Hyper-parameter Value
MetaMetrics-SUM init__\__points 5
n__\__iter 100
MetaMetrics-MT init__\__points 5
n__\__iter 100
MetaMetrics-QA init__\__points 5
n__\__iter 40
MetaMetrics-Cap init__\__points 5
n__\__iter 100
MetaMetrics-RM init__\__points 5
n__\__iter 100
Table 7: Hyper-parameters used for Gaussian Process on each task.
C.2.2 Boosting Method
For XGBoost training, we use a different objective depending on the task. We perform parameter searching with hyper-parameter values in Table 8.
Metric Hyper-parameter Value
MetaMetrics-SUM n__\__estimators__\__low 100
n__\__estimators__\__high 1000
n__\__estimators__\__step 100
n__\__estimators__\__prior Uniform
objective reg::::absoluteerror
MetaMetrics-MT n__\__estimators__\__low 100
n__\__estimators__\__high 1000
n__\__estimators__\__step 100
n__\__estimators__\__prior Uniform
objective reg::::absoluteerror
MetaMetrics-QA n__\__estimators__\__low 100
n__\__estimators__\__high 400
n__\__estimators__\__step 25
n__\__estimators__\__prior Uniform
objective reg::::squaredlogerror
MetaMetrics-Cap n__\__estimators__\__low 100
n__\__estimators__\__high 1000
n__\__estimators__\__step 100
n__\__estimators__\__prior Uniform
objective reg::::squaredlogerror
MetaMetrics-RM n__\__estimators__\__low 100
n__\__estimators__\__high 1000
n__\__estimators__\__step 100
n__\__estimators__\__prior Uniform
objective rank:pairwise
Table 8: Initial hyper-parameter values used for parameter searching during XGBoost.
Metric SummEval Benchmark LLM Combined
Coh. Cons. Fluency Rel. Faith. Coh. Cons. Coh. Cons. Avg.
BLEU 0.110 0.126 0.113 0.170 0.162 0.246 0.252 0.029 0.082 0.143
chrF 0.143 0.094 0.071 0.198 0.272 0.163 0.342 0.068 0.218 0.174
METEOR 0.077 0.102 0.072 0.162 0.251 0.182 0.331 0.033 0.164 0.153
ROUGE-WE1 0.115 0.088 0.081 0.169 0.223 0.218 0.292 0.198 0.272 0.184
ROUGE-1 0.112 0.103 0.061 0.189 0.195 0.252 0.290 0.198 0.248 0.183
ROUGE-3 0.099 0.115 0.048 0.154 0.138 0.142 0.225 0.112 0.191 0.136
ROUGE-4 0.096 0.122 0.067 0.143 0.138 0.085 0.200 0.043 0.148 0.116
ROUGE-l 0.110 0.087 0.069 0.157 0.137 0.270 0.239 0.217 0.210 0.166
BARTScore (Max) 0.084 0.065 0.048 0.135 0.080 0.216 0.166 0.054 0.057 0.101
BARTScore (Mean) 0.086 0.074 0.040 0.143 0.080 0.216 0.166 0.210 0.173 0.132
BLEURT (Max) 0.185 0.070 0.114 0.189 0.046 0.264 0.143 0.229 0.143 0.154
BLEURT (Mean) 0.099 0.011 0.029 0.130 0.046 0.264 0.143 0.265 0.173 0.129
BERTScore (f1) 0.105 0.100 0.120 0.181 0.171 0.280 0.283 0.306 0.304 0.206
BERTScore (Recall) 0.104 0.111 0.093 0.176 0.228 0.223 0.333 0.218 0.300 0.198
SummaQA (F1) 0.051 0.120 0.106 0.114 0.318 0.038 0.293 -0.014 0.180 0.134
SummaQA (Prob) 0.070 0.100 0.070 0.138 0.330 -0.003 0.303 -0.028 0.206 0.132
MetaMetrics-Sum
GP (Full) 0.170 0.142 0.115 0.233 0.356 0.308 0.383 0.315 0.333 0.262
GP (Top 5) 0.149 0.137 0.117 0.211 0.342 0.307 0.370 0.296 0.316 0.249
GP (Oracle Top 5) 0.169 0.138 0.121 0.230 0.356 .308 0.384 0.296 0.316 0.258
XGBoost 0.190 0.151 0.131 0.271 0.406 0.411 0.397 0.403 0.367 0.303
XGBoost (Iterative Best) 0.185 0.165 0.181 0.255 0.406 0.404 0.394 0.402 0.369 0.307
XGBoost (Iterative Top 5) 0.141 0.115 0.160 0.208 0.377 0.378 0.383 0.369 0.351 0.276
XGBoost (Top 5) 0.110 0.096 0.120 0.186 0.314 0.353 0.369 0.343 0.327 0.246
XGBoost (Oracle Top 5) 0.135 0.090 0.134 0.219 0.319 0.354 0.378 0.343 0.325 0.255
Table 9: Test result for summarization task. Coh., Cons., Rel., and Faith. corresponds to coherence, consistency, relevance, and faithfulness respectively. Bold and underlined values indicate the best and second best performance, respectively. The Combined column reflects results when all datasets are merged.
Setup EVOUNA EQAE Combined
NQ TQ NarrativeQA SemEval
ArmoRM Metrics
ultrafeedback-honesty 0.202 0.281 0.297 0.335 0.296
helpsteer-helpfulness 0.204 0.293 0.332 0.427 0.302
helpsteer-correctness 0.206 0.300 0.336 0.403 0.305
argilla-judge-lm 0.243 0.301 0.253 0.337 0.332
n-gram-based Metrics
BLEU1 0.461 0.353 0.362 0.260 0.424
ROUGE 0.494 0.399 0.586 0.521 0.454
ROUGE-L 0.495 0.399 0.584 0.518 0.454
METEOR 0.517 0.435 0.570 0.461 0.494
MetaMetrics-QA
GP (Full) 0.518 0.448 0.586 0.464 0.512
GP (Top 5) 0.506 0.450 0.570 0.452 0.512
GP (Oracle) 0.523 0.446 0.570 0.474 0.512
XGBoost (Full) 0.543 0.471 0.601 0.486 0.536
XGBoost (Top 5) 0.535 0.491 0.624 0.513 0.534
XGBoost (Oracle) 0.543 0.479 0.619 0.520 0.534
XGBoost (Iterative Best) 0.545 0.488 0.626 0.510 0.538
XGBoost (Iterative Top 5) 0.540 0.488 0.626 0.503 0.528
Table 10: Full performance of MetaMetrics-QA compared to the 8 best-performing standalone metrics and other baseline setups. Bold and underlined values indicate the best and second best performance, respectively. The Combined column reflects results when all datasets are merged.
Appendix D Detailed Results
In this section, we provide more detailed results that could not be included in the main paper due to space constraints.
D.1 RewardBench Leaderboard
The leaderboard ranking screenshot is shown in Figure 5.
Figure 5: RewardBench Leaderboard Screenshot (Updated on October 6th, 2024).
Figure 6: RewardBench Detailed Leaderboard Screenshot (Updated on October 6th, 2024).
Metric overall en-de he-en zh-en
sys/seg sys seg seg sys seg seg sys seg seg
avg-corr pearson pearson acc-t pearson pearson acc-t pearson pearson acc-t
Reference-based Metric
    chrF 0.694 0.866 0.232 0.519 0.776 0.221 0.460 0.809 0.063 0.485
    BLEU 0.696 0.917 0.192 0.520 0.769 0.220 0.442 0.734 0.119 0.472
    BERTScore 0.742 0.891 0.325 0.528 0.895 0.335 0.515 0.810 0.236 0.499
    Yisi-1 0.754 0.925 0.366 0.542 0.917 0.395 0.529 0.823 0.290 0.504
    MetricX-23-XXL 0.808 0.977 0.585 0.603 0.910 0.548 0.577 0.873 0.625 0.531
    XCOMET-Ensemble 0.825 0.980 0.695 0.604 0.950 0.556 0.586 0.927 0.650 0.543
    COMET 0.779 0.990 0.432 0.575 0.940 0.401 0.531 0.898 0.396 0.514
    XCOMET-XL 0.817 0.970 0.670 0.596 0.949 0.530 0.576 0.928 0.607 0.530
    XCOMET-XXL 0.817 0.983 0.660 0.602 0.952 0.465 0.560 0.965 0.597 0.546
MetaMetrics-MT
    GP 0.819 0.970 0.638 0.610 0.947 0.546 0.590 0.900 0.646 0.539
    XGBoost 0.825 0.992 0.680 0.616 0.957 0.557 0.574 0.929 0.637 0.546
Reference-free Metric
   mbr-metricx-qe 0.788 0.976 0.571 0.584 0.915 0.411 0.553 0.936 0.489 0.537
   MetricX-23-QE 0.800 0.969 0.626 0.596 0.858 0.520 0.564 0.859 0.647 0.527
   GEMBA-MQM 0.802 0.993 0.502 0.572 0.939 0.401 0.564 0.991 0.449 0.522
   XCOMET-QE-Ensemble 0.808 0.974 0.679 0.588 0.909 0.498 0.554 0.892 0.647 0.533
   CometKiwi-QE 0.781 0.946 0.475 0.569 0.859 0.387 0.544 0.963 0.442 0.525
   MetricX-23-QE-L 0.763 0.868 0.501 0.577 0.616 0.419 0.536 0.835 0.622 0.508
   MetricX-23-QE-XXL 0.797 0.934 0.547 0.607 0.813 0.459 0.575 0.877 0.652 0.531
   CometKiwi-QE-XL 0.786 0.976 0.447 0.571 0.900 0.384 0.533 0.974 0.430 0.522
MetaMetrics-MT-QE
   GP 0.801 0.934 0.556 0.609 0.815 0.474 0.578 0.900 0.660 0.537
   XGBoost 0.805 0.967 0.583 0.604 0.881 0.509 0.568 0.869 0.642 0.526
Table 11: Correlation with human ratings on WMT23 (MQM). The results are collected from Freitag et al. (2023). Bold and underlined values indicate the best and second best performance, respectively.
D.2 Abstractive Text Summarization
Table 9 provides a detailed breakdown of the correlations between various metrics and human preference measurements for abstractive text summarization.
D.3 Machine Translation
Table 11 describes the detailed breakdown of the correlations between metrics with human annotators in WMT23.
D.4 Question Answering
Table 10 provides a detailed breakdown of the correlations between various metrics and human ratings for the Question Answering task.
Metric Flickr8k THumB 1.0 avg.
BLEU-1 0.377 0.264 0.321
BLEU-4 0.361 0.210 0.286
ROUGE 0.375 0.237 0.306
METEOR 0.481 0.281 0.381
CIDEr 0.498 0.263 0.381
SPICE-F 0.591 0.200 0.396
SPICE-PR 0.603 0.170 0.387
SPICE-RE 0.579 0.221 0.400
TIGEr 0.542 0.153 0.348
CLIP-S 0.514 0.292 0.403
RefCLIP-S 0.576 0.210 0.393
MetaMetrics-Cap
GP (Full) 0.664 0.329 0.496
GP (Top 5) 0.655 0.277 0.466
XGBoost (Full) 0.670 0.277 0.474
XGBoost (Top 5) 0.664 0.283 0.474
XGBoost (Iterative-Best) 0.662 0.285 0.474
XGBoost (Iterative-Top5) 0.659 0.276 0.468
MetaMetrics-Cap (Cross-Dataset)
GP (Full) 0.638 0.298 0.468
GP (Top 5) 0.560 0.246 0.403
XGBoost (Full) 0.595 0.323 0.449
XGBoost (Top 5) 0.575 0.311 0.443
XGBoost (Iterative-Best) 0.635 0.321 0.478
XGBoost (Iterative-Top 5) 0.634 0.301 0.467
Table 12: Kendall correlation results of MetaMetrics-Cap is tuned on THumB 1.0 and tested on Flickr8k, and vice versa. Scores are reported for multiple metrics, with the average performance (avg.) across datasets. Bold and underlined values indicate the best and second best performance, respectively.
D.5 Image Captioning
Table 12 provides a detailed breakdown of the correlations between various metrics and human annotations for Image Captioning.
D.6 Reward Model Scoring
Table 13 describes the detailed breakdown of accuracy between metrics with human preference.
Metric Score Chat Chat Hard Safety Reasoning
Llama-3.1-Nemotron-70B-Reward 94.1 97.5 85.7 95.1 98.1
Skywork-Reward-Gemma-2-27B 93.8 95.8 91.4 91.9 96.1
TextEval-Llama3.1-70B 93.5 94.1 90.1 93.2 96.4
Skywork-Critic-Llama-3.1-70B 93.3 96.6 87.9 93.1 95.5
URM-LLaMa-3.1-8B 92.9 95.5 88.2 91.1 97.0
SFR-LLaMa-3.1-70B-Judge-r 92.7 96.9 84.8 91.6 97.6
Skywork-Reward-Llama-3.1-8B 92.5 95.8 87.3 90.8 96.2
GRM-Llama3-8B 91.5 95.5 86.2 90.8 93.6
Internlm2-20B-Reward 90.2 98.9 76.5 89.5 95.8
ArmoRM-Llama3-8B-v0.1 90.4 96.9 76.8 90.5 97.3
URM-LLaMa-3-8B 89.9 96.9 78.7 88.2 95.7
Internlm2-7B-Reward 87.6 99.2 69.5 87.2 94.5
MetaMetrics-RM
GP 93.4 98.3 86.4 90.8 98.2
GP (Top 3) 93.0 95.5 87.7 91.6 97.0
XGBoost 92.8 96.1 88.4 90.9 95.8
XGBoost (Iterative Best) 92.9 95.8 88.4 91.8 95.8
XGBoost (Top 3) 93.0 86.0 90.3 93.0 95.5
Table 13: Accuracy results for Reward-Model-As-A-Metric. Bold and underlined values indicate the best and second best performance, respectively.
Appendix E Weights and Feature Importance
E.1 Question Answering
There are 8 setups for the question answering task. The resulting weights or feature importance for each setup are reported in Figure 7.
Figure 7: Weights and Feature Importance for MetaMetrics-QA.
E.2 Machine Translation
There are 2 setups for MT, reference-based and reference-free. Both setups’ resulting weights or feature importance are reported in 8
(a) Weights and Feature Importance for each reference-based metric used for WMT23 task.
(b) Weights for each reference-free metric used for WMT23 Task.
Figure 8: Weights for MetaMetrics-MT calibrated for WMT MQM task.
E.3 Image Captioning
We utilize two datasets for Image Captioning, each with two setups: a) cross-dataset and b) in-dataset. The resulting weights and feature importance are presented in Figure 9.
Figure 9: Weights and Feature Importance for MetaMetrics-Cap.
E.4 Reward Modeling
The weights and feature importance from the resulting optimization are reported in Figure 9.
Figure 10: Weights and Feature Importance for MetaMetrics-RM.